when a conductor carrying 2.5x10-3 C is discharged to earth, a current of 4mA is needed. How long does the discharge take?


1
Expert's answer
2020-10-21T09:56:35-0400

Solution

Current is given by

"I=\frac{q}{t}"


Time taken to discharge

"t=\frac{q}{I}"

"t=\frac{2.5\times10^{-3}}{4\times10^{-3}}"

"t=0.62sec"

Therefore time taken to discharge is 0.62s.